Based on historical data, your manager believes that 26% of the company's orders come from first-time customers. A random sample of 160 orders will be used to estimate the proportion of first-time-customers. What is the probability that the sample proportion is less than 0.27?

Note: You should carefully round any z-values you calculate to 4 decimal places to match wamap's approach and calculations.

Answer =  (Enter your answer as a number accurate to 4 decimal places.)
Answer #1

Solution

Given that,

\sigma\hat p =  \sqrt[p ( 1 - p ) / n] =  \sqrt [(0.26 * 0.74) / 160 ] = 0.0347

P( \hat p < 0.27) =

= P[( \hat p - \mu \hat p ) / \sigma \hat p < (0.27 - 0.26) / 0.0347]

= P(z < 0.2882)

= 0.6134

Probability = 0.6134